OpenStudy (anonymous):

Two runners start 30 kilometers apart and run toward each other. They meet after 3 hours. One runner runs 2 kilometers per hour faster than the other runner. What rate do they run? A.2 km/hr and 4 km/hr B.3 km/hr and 5 km/hr C.5 km/hr and 7 km/hr D.4 km/hr and 6 km/hr

OpenStudy (tkhunny):

A. 30/(2+4) = 30/6 = 5 -- Too slow! B. 30/(3+5) = 30/8 = 3.75 -- Too slow! C. 30/(5+7) = 30/12 = 2.5 -- Too fast! D, 30/(4+6) = 30/10 = 3 -- Just right.

OpenStudy (anonymous):

Thanks

OpenStudy (tkhunny):

Running directly toward each other, we can just add the rates and cover the distance, ANY other arrangement and it takes more work.

OpenStudy (anonymous):

so the answer is D?